# OpenStereo: A Comprehensive Benchmark for Stereo Matching and Strong Baseline

### v5

@@ Abstract
Stereo matching aims to estimate the disparity between matching pixels in a stereo image pair, which is of great importance to robotics, autonomous driving, and other computer vision tasks. Despite the development of numerous impressive methods in recent years, replicating their results and determining the most suitable architecture for practical application remains challenging. Addressing this gap, our paper introduces a comprehensive benchmark focusing on practical applicability rather than solely on performance enhancement. Specifically, we develop a flexible and efficient stereo matching codebase, called OpenStereo. OpenStereo includes training and inference codes of more than 10 network models, making it, to our knowledge, the most complete stereo matching toolbox available. Based on OpenStereo, we conducted experiments and have achieved or surpassed the performance metrics reported in the original paper. Additionally, we carry out an exhaustive analysis and deconstruction of recent developments in stereo matching through comprehensive ablative experiments. These investigations inspired the creation of StereoBase, a strong baseline model. Our StereoBase ranks 1st on SceneFlow, KITTI 2015, 2012 (Reflective) among published methods and achieves the best performance across all metrics. In addition, StereoBase has strong cross-dataset generalization. Code is available at https://github.com/XiandaGuo/OpenStereo.
11footnotetext: Joint second authors.44footnotetext: Corresponding authors
Figure 1: Comparison with other state-of-the-art stereo methods [24, 40, 1, 46, 5, 13, 39, 12, 51] on SceneFlow [28], KITTI 2012 [11], 2015 [29] leaderboards.
@@ 1 Introduction
Traditional stereo matching algorithms typically rely on matching corresponding image regions between the left and right views based on their similarity measures. Several techniques have been proposed in the literature for stereo matching, including methods based on gray-level information [2, 23, 56], region-based approaches [62, 33], and energy optimization methods [37, 16]. Gray-level-based methods [2, 23, 56] compute disparities by matching pixels or groups of pixels with similar intensities in the left and right images. Region-based approaches [62, 33] group pixels into larger regions and then compute correspondences between these regions. Energy optimization methods [37, 16] formulate the stereo matching problem as an energy minimization problem, where the goal is to find the disparity map that minimizes a certain energy function. While these methods have achieved some success in stereo matching, they suffer from certain limitations and challenges, such as occlusions, textureless regions, and computational complexity.
Recently, with the support of large synthetic datasets [28, 36, 38, 11, 29], CNNs-based stereo matching methods [19, 5, 58, 13, 52, 25, 7, 46] has achieved impressive results. These methods have shown remarkable performance in terms of accuracy and efficiency. Unlike traditional methods, CNN-based stereo matching methods are end-to-end frameworks that take a pair of stereo images as input and directly predict the disparity between them. In addition to the accuracy and efficiency benefits, CNN-based stereo matching methods have shown the potential for generalization to new environments and scenes beyond the training data. This is mainly due to the ability of CNNs to learn feature representations from raw data that are invariant to scale, rotation, and illumination changes. Furthermore, the availability of large-scale datasets has enabled the use of deep learning techniques for stereo matching, allowing for more complex and sophisticated models to be trained.
Nevertheless, we find that not all datasets are accompanied by official evaluation tools. For example, the DrivingStereo [55] dataset does not provide specific evaluation scripts, making comparative assessments challenging. The SceneFlow [28] dataset, with its finalpass and cleanpass data varieties, complicates fair model comparisons. Generalization experiments for stereo matching algorithms typically train on the SceneFlow dataset and evaluate on KITTI2012 [11], KITTI2015 [29], ETH3D [38], and Middlebury [36]. However, the inconsistency in evaluation data volume and metric definitions across studies leads to comparative discrepancies. Furthermore, even when source codes for models are available, reproducing the reported benchmarks is exceedingly challenging, e.g. STTR [24]. Hence, there’s a pressing need for a comprehensive benchmark study within the stereo matching community to enhance practicality and ensure consistent comparisons. To achieve this objective, we introduce a versatile stereo matching codebase: OpenStereo.
@@
To promote scalability and adaptability, OpenStereo offers the following features: (1) Modual design, researchers can define a new model without the need to alter the model code itself by simply modifying a yaml configuration file. (2)Various frameworks, including Concatenation-based [5, 58], Correlation-based [52, 45, 46, 1], Interlaced-based [39], Group-wise-correlation-based [51], Combine-based methods [13], and Difference-based [21]. (3)Various datasets, including SceneFlow [28], KITTI2012 [11], KITTI2015 [29], Middlebury [36], ETH3D [38] and DrivingStereo [55] dataset. (4) State-of-the-art methods, including PSMNet [5], GwcNet [13], AANet [52], FADNet++ [46], CFNet [40], STTR [24], CoEx [1], CascadeStereo [12], MobileStereoNet [39] and IGEV [51].
@@
Leveraging OpenStereo, we rigorously reassess various officially stated conclusions by meticulously re-implementing the ablation studies, including data augmentation, backbone architectures, cost construction, disparity regression and refinement processes. Based on the insights gleaned from these ablation experiments, we introduce StereoBase, a model that sets a new benchmark, surpassing recently proposed methods in terms of performance. Furthermore, we undertake a thorough investigation to establish that StereoBase is not only powerful but also serves as an empirically robust baseline model for stereo matching, demonstrating exceptional efficacy and resilience across diverse testing scenarios.
In summary, this paper makes significant contributions to the field of stereo matching research from three key aspects. (1) OpenStereo, a unified and extensible platform, enables researchers to conduct comprehensive stereo matching studies. (2) With the support of OpenStereo, we conduct a profound revisitation and thorough deconstruction of recent stereo matching methodologies. (3) We introduce StereoBase, which sets a new benchmark with EPE of 0.34 on SceneFlow [28] and ranks 1st on KITTI2015 [29] and 2012(Reflective)[11] leaderboards among published methods. Furthermore, StereoBase exhibits exceptional cross-dataset generalization capabilities. These contributions collectively advance the state-of-the-art in stereo matching, offering both insights and practical tools to the stereo matching research community.
@@ 2.1 Stereo Matching
Figure 2: Timeline of Stereo Matching Models. This figure illustrates the timeline of various stereo matching models that have been proposed in the literature. The top part shows ED-conv2D-based models, while the bottom part shows CVM-conv3D-based models. Each model is labeled with its name and authors.
Stereo Matching with CVM-Conv3D The CVM-Conv3D methods are proposed to improve the performance of depth estimation [19, 5, 13, 58, 59, 12, 7, 44, 1, 41, 39, 22, 26, 50, 31]. These methods learn disparities from a 4D cost volume, which is mainly constructed by concatenating left feature maps with their corresponding right counterparts across each disparity level [5]. GCNet [19] firstly introduced a novel approach that combines 3D encoder-decoder architecture with a 2D convolutional network to obtain a dense feature representation, which is used to regularize a 4D concatenation volume. Following GCNet [19], PSMNet [5] proposes an approach for regularizing the concatenation volume by leveraging a stacked hourglass 3D convolutional neural network in tandem with intermediate supervision. In order to enhance the expressiveness of the cost volume and ultimately improve performance in ambiguous regions, GwcNet [58] proposes the group-wise correlation volume and ACVNet [50] proposes the attention concatenation volume. GANet [58] proposes a semi-global aggregation layer and local guided aggregation layer to capture local and the whole-image cost dependencies respectively. Based on GANet [58], DSMNet [59] construct a robust “domain normalization” method to overcome the challenges in cross-domain generalization. CoEx [1] proposes a novel approach called Guided Cost volume Excitation (GCE), which leverages image guidance to construct a simple channel excitation of the cost volume. IGEV-Stereo [51] leverages an iterative geometry encoding volume to capture both local and non-local geometry information, which outperforms existing methods on KITTI benchmarks and achieves cross-dataset generalization and high inference efficiency. Besides, as a potential solution to reduce the need for humans to design a good volumetric method, Neural Architecture Search (NAS) [15, 27, 3] has been employed in stereo matching [7].
@@
Stereo Matching with ED-Conv2D The ED-Conv2D methods [57, 28, 21, 34, 54, 9, 35, 46, 52, 43, 40, 24, 25, 39, 61], which adopt networks with 2D convolutions to predict disparity, has been driven by the need for improved accuracy, computational efficiency, and real-time performance. One of the early deep learning-based stereo matching methods, MC-CNN (Matching Cost CNN) [57], was proposed to learn a matching cost function for improving performance in the cost aggregation and optimization stages. Subsequently, Mayer et al. [28] present end-to-end networks for the estimation of disparity, called DispNet, which is pure 2D CNN architectures. However, the model still faces challenges in capturing the matching features, resulting in poor estimation results. To overcome this challenge, the correlation layer is introduced in the end-to-end architecture [28, 8, 17, 18] to better capture the relationship between the left and right images. By incorporating this layer, the accuracy of the model is significantly improved. Furthermore, FADNet++ [46] proposes an innovative approach to efficient disparity refinement using residual learning [14] in a coarse-to-fine manner. AutoDispNet [35] applied neural architecture search to automatically design stereo matching network structures. More recently, Croco-Stereo [47] shows that large-scale pre-training can be successful for stereo matching through well-adapted pretext tasks. This method can achieve state-of-the-art performance without using task-specific designs, like correlation volume or iterative estimation.
@@ 3.1 Design Principles of OpenStereo
Support for various evaluation datasets. OpenStereo is a comprehensive tool that includes datasets commonly used by researchers in stereo matching. OpenStereo not only supports synthetic stereo datasets such as SceneFlow [28], but also five real-world datasets: KITTI2012 [11], KITTI2015 [29], ETH3D [38], Middlebury [36], and DrivingStereo [55]. We introduce a suite of bespoke functions, meticulously crafted for each dataset, encompassing everything from initial data preprocessing to the final stages of evaluation. The evaluation module includes the submission of the results to all benchmarks.
Support for state-of-the-arts. In our work, we have successfully replicated various state-of-the-art stereo matching methods, including PSMNet [5], GwcNet [13], AANet [52], FADNet++ [46], CFNet [40], STTR [24], CoEx [1], CascadeStereo [12], MobileStereoNet [39] and IGEV [51]. Notably, the performance metrics we achieved, in most cases, surpass those reported in their original publications. This comprehensive replication effort provides a valuable resource for novices in the field, offering a rich array of official examples to facilitate an effective and efficient start in stereo matching research.
@@ 4.1 Recheck on Previous Methods
PSMNet [5] is one of the most crucial and frequently cited networks in stereo matching. The official implementation code is available at https://github.com/pytorch/vision. GwcNet [13] employs a Group-wise Correlation Stereo Network to exploit both local and global matching cost volumes (https://github.com/xy-guo/GwcNet). AANet [52] focuses on adaptively fusing multi-scale features through a lightweight and efficient adaptive aggregation module ( https://github.com/haofeixu/aanet). FADNet++ [46] employs a two-stage architecture, consisting of a feature-metric aggregation stage and a disparity refinement stage. (https://github.com/HKBU-HPML/FADNet). CFNet [40] utilizes a collaborative feature learning strategy to enhance the representation capacity of features in the matching cost computation (https://github.com/gallenszl/CFNet). STTR [24] focuses on geometric correspondence problems, including stereo matching, by employing transformers to capture long-range spatial and temporal context information (https://github.com/mli0603/stereo-transformer). CoEx [1] focuses on context-aware features and cost volume aggregation (https://github.com/antabangun/coex). CascadeStereo [12] adopts a cascaded architecture for accurate and efficient disparity estimation, whose official code is available at (https://github.com/alibaba/cascade-stereo). MobileStereoNet [39] builds upon the StereoNet [21], which proposes Interlacing Cost Volume Construction (https://github.com/cogsys-tuebingen/mobilestereonet). IGEV [51] builds a combined geometry encoding volume that not only encodes both geometric and contextual information but also intricately captures fine-grained details of local matching (https://github.com/gangweiX/IGEV).
Table 1: Quantitative evaluation on the SceneFlow [28] test. For each model, the specific category on the SceneFlow used is consistent with the original paper. Models marked with are OpenStereo implementations, others are from the original paper. Underline refers to evaluation in the non-occluded regions only [24]. Bold: Best
Method SceneFlow
TrainSize Categories EPE(px)
STTR [24] $256\times 512256 × 512$ FinalPass 0.43
STTR [24] $256\times 512256 × 512$ FinalPass 0.40
PSMNet [5] $256\times 512256 × 512$ CleanPass 1.09
PSMNet [5] $256\times 512256 × 512$ CleanPass 0.96
CFNet [40] $256\times 512256 × 512$ FinalPass 1.04
CFNet [40] $256\times 512256 × 512$ FinalPass 0.96
AANet [52] $288\times 576288 × 576$ CleanPass 0.87
AANet [52] $288\times 576288 × 576$ CleanPass 0.80
Mobilestereo-2D [39] $256\times 512256 × 512$ FinalPass 1.14
Mobilestereo-2D [39] $256\times 512256 × 512$ FinalPass 1.10
Mobilestereo-3D [39] $256\times 512256 × 512$ FinalPass 0.80
Mobilestereo-3D [39] $256\times 512256 × 512$ FinalPass 0.79
GwcNet [13] $256\times 512256 × 512$ FinalPass 0.76
GwcNet [13] $256\times 512256 × 512$ FinalPass 0.74
COEX [1] $288\times 576288 × 576$ FinalPass 0.68
COEX [1] $288\times 576288 × 576$ FinalPass 0.67
FADNet++ [46] $384\times 768384 × 768$ CleanPass 0.76
FADNet++ [46] $384\times 768384 × 768$ CleanPass 0.65
CascadeStereo [12] $256\times 512256 × 512$ FinalPass 0.72
CascadeStereo [12] $256\times 512256 × 512$ FinalPass 0.67
IGEV [51] $320\times 736320 × 736$ FinalPass 0.47
IGEV [51] $320\times 736320 × 736$ FinalPass 0.46
Table 2: Quantitative evaluation on KITTI2015 leadboard.
Method KITTI2015
D1-bg D1-fg D1-all
PSMNet [5] 1.86 4.62 2.32
PSMNet [5] 1.80 4.58 2.26
IGEV [51] 1.38 2.67 1.59
IGEV [51] 1.44 2.31 1.59
For benchmarking, it is critical to ensure that the results are reliable and trustworthy. To achieve this, we conducted our experiments using SceneFlow [28] and KITTI2015 [29] datasets. As shown in Table 1, in most cases, the reproduced performances of OpenStero are better than the results reported by the original papers. Regarding the KITTI2015 dataset, submission constraints led us to limit our leaderboard contributions to reproductions of the widely recognized PSMNet [5] and the latest state-of-the-art IGEV [51], as demonstrated in Table 2. We stand by the integrity of our benchmark, asserting its role in providing a reliable and valid evaluation of stereo matching models. OpenStereo is designed to offer the research community in stereo matching a standardized, comprehensive platform for method assessment. This facility enables meaningful and comparative analyses across various models, thereby serving as an valuable tool for the innovation and scrutiny of emerging algorithms.
@@ 4.2 Denoising Stereo Matching
With the support of OpenStereo, a comprehensive reevaluation of various stereo matching methods is conducted, including data augmentation, feature extraction, cost construction, disparity prediction and refinement.
Table 3: The model undergoes training and evaluation on the finalpass of the SceneFlow [28] training and test sets. KITTI2015 [29] training set, consisting of 200 images, is only employed to evaluate the generalizability of models under various data augmentation conditions. RC stands for RandomCrop. CES represents ColorAug, Erase, and Scale. HFlip denotes both images of a stereo and disparity are horizontally flipped. HSFlip horizontally flips both images in the stereo pair and then swaps them. VFlip involves vertically flipping both images in the stereo pair along with the disparity, inverting their top-bottom orientation. denotes the method proposed in this paper.
Data Augmentation SceneFlow KITTI15
@@
4.2.1 Data Augmentation
While five data augmentation techniques are prevalent in stereo matching [25, 51], including random crop, color augmentation, eraser transform, flip, and spatial transform, the empirical efficacy of each method specifically for stereo vision tasks remains an underexplored area in the field. Our work extends beyond these common practices, exploring not only these techniques but also experimenting with two additional methods: BatchRandom and Shift, to further understand their impact on stereo matching performance. BatchRandom denotes the technique applied during model training where the size of the random crop varies for each batch. The Shift involves randomly shifting the right image horizontally by a certain range and then cropping both images to a specified size. As shown in Table 3, most data augmentation strategies, except for the combined use of erase transform and random crop, lead to a decline in the model’s performance metrics of SceneFlow. This is because stereo matching involves pixel-level matching, and these data augmentations (color augmentation, flip, and spatial transform) can affect the alignment of pixels. The erase transform benefits stereo matching by enhancing the model’s capacity to cope with occlusions, missing data, and diverse real-world scenarios. It encourages robust feature learning and generalization, ultimately improving model performance. Furthermore, the combination of RandomCrop(320×736) with color augmentation delivers the most striking improvements in Kitti2015 dataset performance, achieving the lowest EPE of 1.61 px and the lowest D1_all of 7.81. Such improvements highlight the critical role of color augmentation in boosting the model’s resilience against the variability of color and lighting that naturally occurs in diverse driving environments.
Table 4: Ablation study of different backbones and pre-training methods on the finalpass of SceneFlow test datasets [28]. Flops and Params represent the computational complexity and parameters within the whole model, respectively.
Backbone Pre-train Flops Params EPE
FadNet++ [46] None 28.39G 1.77M 0.8633
MobilenetV2 100 [51] None 31.83G 2.78M 0.8227
PSMNet [5] None 130.87G 4.46M 0.8068
MobilenetV2 100 [51] None 31.83G 2.78M 0.8227
MobilenetV2 100 [51] Timm [48] 31.83G 2.78M 0.7132
MobilenetV2 120d [51] Timm [48] 38.75G 5.21M 0.6603

### v6

@@ Abstract
Stereo matching aims to estimate the disparity between matching pixels in a stereo image pair, which is important to robotics, autonomous driving, and other computer vision tasks. Despite the development of numerous impressive methods in recent years, determining the most suitable architecture for practical application remains challenging. Addressing this gap, our paper introduces a comprehensive benchmark focusing on practical applicability rather than solely on individual models for optimized performance. Specifically, we develop a flexible and efficient stereo matching codebase, called OpenStereo. OpenStereo includes training and inference codes of more than 10 network models, making it, to our knowledge, the most complete stereo matching toolbox available. Based on OpenStereo, we conducted experiments and have achieved or surpassed the performance metrics reported in the original paper. Additionally, we conduct an exhaustive analysis and deconstruction of recent developments in stereo matching through comprehensive ablative experiments. These investigations inspired the creation of StereoBase, a strong baseline model. Our StereoBase ranks 1st on SceneFlow, KITTI 2015, 2012 (Reflective) among published methods and performs best across all metrics. In addition, StereoBase has strong cross-dataset generalization. Code is available at https://github.com/XiandaGuo/OpenStereo.
11footnotetext: Joint second authors.22footnotetext: Corresponding authors:long.chen@ia.ac.cn
@@ 1 Introduction
Traditional stereo-matching algorithms typically rely on matching corresponding image regions between the left and right views based on their similarity measures. Several techniques have been proposed in the literature for stereo matching, including methods based on gray-level information [3, 29, 68], region-based approaches [73, 39], and energy optimization methods [44, 21]. With the support of large synthetic datasets [35, 43, 45, 14, 36, 66], CNNs-based stereo matching methods [24, 6, 70, 18, 61, 33, 10, 52, 60, 53, 8, 30, 32] has achieved impressive results. As shown in Figure 1, based on the network pipeline of stereo matching, CNNs-based stereo matching methods can be roughly grouped into two categories [52], including the encoder-decoder network with 2D convolution (ED-Conv2D) [35, 40, 65, 41, 51, 52, 61, 49, 31, 33, 46, 28, 56, 30, 16] and the cost volume matching with 3D convolution (CVM-Conv3D) [24, 6, 67, 55, 70, 18, 71, 12, 74, 15, 1, 10, 2, 47, 59, 60, 54, 9, 62].
However, we find that different studies often employ varying data augmentation strategies, learning rates, learning rate optimization methods, and backbone architectures. This inconsistency makes it difficult to evaluate and compare various methods’ performance accurately. This inconsistency in experimental setups and methodologies makes it difficult to derive conclusive insights and hampers the objective assessment of advancements in stereo matching. Without a standardized benchmark, researchers struggle to identify the true impact of individual components and innovations. There is a lack of clear conclusions and exploration regarding data augmentation strategies, backbone selection, and cost construction methods in stereo matching.
Moreover, not all datasets are accompanied by official evaluation tools. For example, the DrivingStereo [66] dataset does not provide specific evaluation scripts, making comparative assessments challenging. The SceneFlow [35] dataset, with its finalpass and cleanpass data varieties, complicates fair model comparisons. Generalization experiments for stereo matching algorithms typically train on the SceneFlow dataset and evaluate on KITTI2012 [14], KITTI2015 [36], ETH3D [45], and Middlebury [43]. However, due to the absence of a standard protocol for generalization experiments, different papers may yield inconsistent results for the same method. For instance, discussions on the generalization performance of IGEV [60] across works [60, 53, 17] exemplify this issue.
Figure 1: Timeline of Stereo Matching Models. The top part shows ED-conv2D-based models, while the bottom part shows CVM-conv3D-based models. Each model is labeled with its name and authors.
@@
Hence, there’s a pressing need for a comprehensive benchmark study within the stereo-matching community to enhance practicality and ensure consistent comparisons. To achieve this objective, we introduce a versatile stereo-matching codebase: OpenStereo. To promote scalability and adaptability, OpenStereo offers the following features: (1) Modual design, researchers can define a new model without the need to alter the model code itself by simply modifying a yaml configuration file. (2)Various frameworks, including Concatenation-based [6, 70], Correlation-based [61, 51, 52, 2], Interlaced-based [46], Group-wise-correlation-based [60], Combine-based methods [18], and Difference-based [26]. (3)Various datasets, including SceneFlow [35], KITTI2012 [14], KITTI2015 [36], Middlebury [43], ETH3D [45] and DrivingStereo [66] dataset. (4) State-of-the-art methods, including PSMNet [6], GwcNet [18], AANet [61], FADNet++ [52], CFNet [47], STTR [31], CoEx [2], CascadeStereo [15], MobileStereoNet [46] and IGEV [60].
@@
Leveraging OpenStereo, we rigorously reassess various officially stated conclusions by meticulously re-implementing the ablation studies, including data augmentation, backbone architectures, cost construction, disparity regression, and refinement processes. Based on the insights gleaned from these ablation experiments, we introduce StereoBase, a model that sets a new benchmark, surpassing recently proposed methods in terms of performance. StereoBase is powerful and serves as an empirically robust baseline model for stereo matching, demonstrating exceptional efficacy and resilience across diverse testing scenarios.
Our contributions are summarized as follows:
  • $\bullet$
    We introduce OpenStereo, a unified and extensible platform, which enables researchers to conduct comprehensive stereo matching studies.
  • $\bullet$
    We conduct a profound revisitation and thorough deconstruction of recent stereo-matching methodologies.
  • $\bullet$
    We introduce StereoBase, which sets a new benchmark with EPE of 0.34 on SceneFlow [35] and ranks 1st on KITTI2015 [36] and 2012(Reflective)[14] leaderboards among published methods.
@@ 2.1 Stereo Matching
Stereo Matching with CVM-Conv3D The CVM-Conv3D methods are proposed to improve the performance of depth estimation [24, 6, 67, 55, 70, 18, 71, 12, 74, 15, 1, 10, 2, 47, 59, 60, 54, 9, 62]. These methods learn disparities from a 4D cost volume, mainly constructed by concatenating left feature maps with their corresponding right counterparts across each disparity level [6]. GCNet [24] firstly introduced a novel approach that combines 3D encoder-decoder architecture with a 2D convolutional network to obtain a dense feature representation, which is used to regularize a 4D concatenation volume. Following GCNet [24], PSMNet [6] proposes an approach for regularizing the concatenation volume by leveraging a stacked hourglass 3D convolutional neural network in tandem with intermediate supervision. To enhance the expressiveness of the cost volume and ultimately improve performance in ambiguous regions, GwcNet [70] proposes the group-wise correlation volume and ACVNet [59] proposes the attention concatenation volume. GANet [70] proposes a semi-global aggregation layer and local guided aggregation layer to capture local and the whole-image cost dependencies respectively. Based on GANet [70], DSMNet [71] constructs a robust “domain normalization” method to overcome the challenges in cross-domain generalization. CoEx [2] proposes a novel approach called Guided Cost volume Excitation (GCE), which leverages image guidance to construct a simple channel excitation of the cost volume. IGEV-Stereo [60] leverages an iterative geometry encoding volume to capture local and non-local geometry information, outperforming existing methods on KITTI benchmarks and achieving cross-dataset generalization and high inference efficiency. Besides, as a potential solution to reduce the need for humans to design a good volumetric method, Neural Architecture Search (NAS) [20, 34, 4] has been employed in stereo matching [10].
@@
Stereo Matching with ED-Conv2D The ED-Conv2D methods [35, 40, 65, 41, 51, 52, 61, 49, 31, 33, 46, 28, 56, 30, 16], which adopt networks with 2D convolutions to predict disparity, has been driven by the need for improved accuracy, computational efficiency, and real-time performance. One of the early deep learning-based stereo matching methods, MC-CNN (Matching Cost CNN) [69], was proposed to learn a matching cost function for improving performance in the cost aggregation and optimization stages. Subsequently, Mayer et al [35] present end-to-end networks for the estimation of disparity, called DispNet, which is pure 2D CNN architectures. However, the model still faces challenges in capturing the matching features, resulting in poor estimation results. To overcome this challenge, the correlation layer is introduced in the end-to-end architecture [35, 11, 22, 23] to better capture the relationship between the left and right images. By incorporating this layer, the accuracy of the model is significantly improved. Furthermore, FADNet++ [52] proposes an innovative approach to efficient disparity refinement using residual learning [19] in a coarse-to-fine manner. AutoDispNet [41] applied neural architecture search to automatically design stereo matching network structures. More recently, Croco-Stereo [56] shows that large-scale pre-training can be successful for stereo matching through well-adapted pretext tasks. This method can achieve state-of-the-art performance without using task-specific designs, like correlation volume or iterative estimation.
@@ 3.1 Design Principles of OpenStereo
Support for various evaluation datasets. OpenStereo is a comprehensive tool that not only supports synthetic stereo datasets such as SceneFlow [35], but also five real-world datasets: KITTI2012 [14], KITTI2015 [36], ETH3D [45], Middlebury [43], and DrivingStereo [66]. We introduce a suite of bespoke functions, meticulously crafted for each dataset, encompassing everything from initial data preprocessing to the final stages of evaluation. The evaluation module includes the submission of the results to KITTI2012 [14] and KITTI2015 [36] leadboards.
Support for state-of-the-arts. In our work, we have successfully replicated various state-of-the-art stereo matching methods, including PSMNet [6], GwcNet [18], AANet [61], FADNet++ [52], CFNet [47], STTR [31], CoEx [2], CascadeStereo [15], MobileStereoNet [46] and IGEV [60]. As shown in Figure 3, the performance metrics we achieved, in most cases, surpass those reported in their original publications.
@@ 4.1 Recheck on Previous Methods
Figure 3: Quantitative evaluation on the SceneFlow [35] test and KITTI2015 [36] leadboard. For each model, the specific category on the SceneFlow used is consistent with the original paper. Underline refers to evaluation in the non-occluded regions only STTR [31].
For benchmarking, it is critical to ensure that the results are reliable and trustworthy. To achieve this, we conducted our experiments on SceneFlow [35] and KITTI2015 [36] datasets. As shown in Figure 3, the reproduced performances of OpenStereo are better than the results reported by the original papers. Regarding the KITTI2015 dataset, submission constraints led us to limit our leaderboard contributions to reproductions of the widely recognized PSMNet [6] and the latest state-of-the-art IGEV [60]. OpenStereo is designed to offer the research community in stereo matching a standardized, comprehensive platform for method assessment. This facility enables meaningful and comparative analyses across various models, thereby serving as an valuable tool for the innovation and scrutiny of emerging algorithms.
@@ 4.2 Denoising Stereo Matching
With the support of OpenStereo, a comprehensive reevaluation of various stereo-matching methods is conducted, including data augmentation, feature extraction, cost construction, disparity prediction, and refinement. Our ablation studies have revealed some new insights.
Figure 4: The visualization of stereo images and disparity with different data augmentation. The blue box represents the area where random cropping occurs during training. Notably, when the views are horizontally flipped, the disparity are multiplied by 11-1- 1, making the disparity map appear pure white.
Table 1: The model undergoes training and evaluation on the finalpass of the SceneFlow [35] training and test sets. KITTI2015 [36] training set, consisting of 200 images, is only employed to evaluate the generalizability of models. RC stands for RandomCrop. HFlip denotes both images of a stereo and disparity are horizontally flipped. HSFlip horizontally flips both images in the stereo pair and then swaps them. VFlip involves vertically flipping both images in the stereo pair along with the disparity, inverting their top-bottom orientation. CES represents ColorAug, Erase, and Scale. Settings used in our final model are underlined.
Data Augmentation LR_scheduler SceneFlow KITTI15
@@
[…]
Table 2: Ablation study of different backbones on the finalpass of SceneFlow test datasets [35]. Flops and Params represent the computational complexity and parameters within the whole model, respectively. FLOPs are calculated at a resolution of $544\times 960544 × 960$.
Backbone Type Pretrain Flops Params EPE
MobilenetV2 100 [42] CNN 70.58G 2.78M 0.7737
MobilenetV2 100 [42] CNN 70.58G 2.78M 0.6155
MobilenetV2 100 [42] CNN 70.58G 2.78M 0.6155
MobilenetV2 120d [42] CNN 85.93G 5.21M 0.5573
EfficientNetV2 [48] CNN 157.52G 24.92M 0.5207
RepViT [50] Trans. 101.35G 5.64M 0.5858
MPViT [27] CNN&Trans. 283.35G 13.33M 0.5113
4.2.1 LR_scheduler and Data Augmentation
As shown in Table 1, MultiStepLR yields an EPE of 0.6839, while OneCycleLR achieves a lower EPE of 0.6155. This substantial difference underscores the crucial role of selecting an appropriate learning rate scheduler for stereo matching. The superior performance of OneCycleLR indicates its potential to improve model accuracy and robustness, making it a preferable choice over MultiStepLR for training stereo-matching models. Although five data augmentation techniques—random crop, color augmentation, eraser transform, flip, and spatial transform—are commonly used in stereo matching [33, 60], their empirical efficacy specifically for stereo matching has not been thoroughly explored. This study investigated these data augmentation strategies to address this gap. Visualizations of these augmentations are provided in Figure 4 for further insight. As shown in Table 1, most data augmentation strategies, except for the combined use of HSFlip and random crop, lead to a decline in the model’s EPE metric of SceneFlow. This is because stereo matching involves pixel-level matching, and these data augmentations (color augmentation and spatial transform) can affect the alignment of pixels. The combination of ColorAug, Erase, and Scale (CES) shows the best generalization performance on KITTI2015, with the lowest EPE of 1.56 and D1_all of 7.64, although it increases the EPE on SceneFlow to 0.7240. These findings underscore the importance of selecting appropriate data augmentation methods to enhance model accuracy and robustness.